2 項分布とカイ 2 乗分布

平成 20 年 3 月 20 日

1 カイ 2 乗分布

これはカイ 2 乗分布のパラメータ ν = 10, 7, 5, χ(10), χ(7), χ(5) がつぎで与えられます。

図 1:

f(x) = (const)x^{ν/2−1} exp(−x/2), x > 0

Plot[{PDF[ChiSquareDistribution[10], x],
PDF[ChiSquareDistribution[7], x],
PDF[ChiSquareDistribution[5], x]},
{x, 0, 25}]
\begin{center}
\includegraphics[width=5cm, height=5cm]
  {ChiSquare.eps}
\end{center}